Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
9.5
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
9.5

Dr. Jao worked on the issue of the NOS/DB status. He observed the glitch and found is enough to click UPDATe on the NOS webpage instead of log-out/log-in. He also observed than, even under the glitch, the DB is correctly updated. Then, maybe this is only a local issue, on the CTIMac3. Thanks a lot to Wei-Chun for his time, patience and help!
Did 4 tries on GAI1859-5855_gapZ, beyond limits! lost track/pointing. It took me time to recover, telescope did not move and readouts showed telescope falling to the EAST! TCS screen reported message DRV DIS (???) Please, could TELOPS, mechanics check, just in case? Thanks in advance.
Finally parked.